\begin{thebibliography}{99}

\bibitem{Beckner08}
W. Beckner,
Weighted inequalities and Stein-Weiss potentials.
\emph{Forum Math.} \textbf{20} (2008), no. 4, 587--606.


\bibitem{Brezis18}
H. Brezis, P. Mironescu,
Gagliardo-Nirenberg inequalities and non-inequalities: the full story.
\emph{Ann. Inst. H. Poincar\'{e} Anal. Non Lin\'{e}aire} \textbf{35} (2018), no. 5, 1355--1376.


\bibitem{Caffarelli84}
L. Caffarelli, R. Kohn, L. Nirenberg,
First order interpolation inequalities with weights. \emph{Compositio Math.} \textbf{53} (1984), no. 3, 259--275.


\bibitem{Catrina09}
F. Catrina, D.G. Costa,
Sharp weighted-norm inequalities for functions with compact support in $\mathbb{R}^N\setminus\{0\}$.
\emph{J. Differential Equations} \textbf{246} (2009), no. 1, 164--182.


\bibitem{Catrina01}
F. Catrina, Z.-Q. Wang,
On the Caffarelli-Kohn-Nirenberg inequalities: sharp constants, existence (and nonexistence), and symmetry of extremal functions.
\emph{Comm. Pure Appl. Math.} \textbf{54} (2001), no. 2, 229--258.


\bibitem{Cazacu20}
C. Cazacu,
A new proof of the Hardy-Rellich inequality in any dimension.
\emph{Proc. Roy. Soc. Edinburgh Sect. A} \textbf{150} (2020), no. 6, 2894--2904.


\bibitem{Cazacu21}
C. Cazacu, J. Flynn, N. Lam,
Short proofs of refined sharp Caffarelli-Kohn-Nirenberg inequalities.
\emph{J. Differential Equations} \textbf{302} (2021), 533--549.


\bibitem{Cazacu22}
C. Cazacu, J. Flynn, N. Lam,
Sharp second order uncertainty principles.
\emph{J. Funct. Anal.} \textbf{283} (2022), no. 10, Paper No. 109659, 37 pp.


\bibitem{Cazacu23}
C. Cazacu, J. Flynn, N. Lam,
Caffarelli-Kohn-Nirenberg inequalities for curl-free vector fields and second order derivatives.
\emph{Calc. Var. Partial Differential Equations} \textbf{62} (2023), no. 4, Paper No. 118, 26 pp.


\bibitem{Cazacu24-2}
C. Cazacu, J. Flynn, N. Lam, G. Lu,
Caffarelli-Kohn-Nirenberg identities, inequalities and
their stabilities.
\emph{J. Math. Pures Appl. (9)} \textbf{182} (2024), 253--284.


\bibitem{Cordero-Erausquin04}
D. Cordero-Erausquin, B. Nazaret, C. Villani,
A mass-transportation approach to sharp Sobolev and Gagliardo-Nirenberg inequalities.
\emph{Adv. Math.} \textbf{182} (2004), no. 2, 307--332.


\bibitem{Costa08}
D.G. Costa,
Some new and short proofs for a class of Caffarelli-Kohn-Nirenberg type inequalities.
\emph{J. Math. Anal. Appl.} \textbf{337} (2008), no. 1, 311--317.


\bibitem{DelPino02}
M. Del Pino, J. Dolbeault,
Best constants for Gagliardo-Nirenberg inequalities and applications to nonlinear diffusions.
\emph{J. Math. Pures Appl. (9)} \textbf{81} (2002), no. 9, 847--875.


\bibitem{DelPino03}
M. Del Pino, J. Dolbeault,
The optimal Euclidean $L^p$-Sobolev logarithmic inequality.
\emph{J. Funct. Anal.} \textbf{197} (2003), no. 1, 151--161.


\bibitem{Do23}
A.X. Do, J. Flynn, N. Lam, G. Lu,
$L^{p}$-Caffarelli-Kohn-Nirenberg inequalities and their stabilities. arXiv:2310.07083.
\url{https://arxiv.org/abs/2310.07083}


\bibitem{Dong18}
M. Dong, N. Lam, G. Lu,
Sharp weighted Trudinger-Moser and Caffarelli-Kohn-Nirenberg inequalities and their extremal functions.
\emph{Nonlinear Anal.} \textbf{173} (2018), 75--98.


\bibitem{Duong21}
A.T. Duong, V.H. Nguyen,
The sharp second order Caffareli-Kohn-Nirenberg inequality and stability estimates for the sharp second order uncertainty principle. arXiv:2102.01425. \url{https://arxiv.org/abs/2102.01425}


\bibitem{Duong25}
A.T. Duong, V.H. Nguyen,
On the sharp second order Caffarelli-Kohn-Nirenberg inequality.
\emph{Ann. Fenn. Math.} \textbf{50} (2025), no. 1, 275--286.


\bibitem{Duy22}
N.T. Duy, N. Lam, G. Lu,
$p$-Bessel pairs, Hardy's identities and inequalities and Hardy-Sobolev inequalities with monomial weights.
\emph{J. Geom. Anal.} \textbf{32} (2022), no. 4, Paper No. 109, 36 pp.


\bibitem{Frank11}
R.L. Frank,
Sobolev inequalities and uncertainty principles in mathematical physics: Part I. Lecture Notes. (2011)
\url{https://www.mathematik.uni-muenchen.de/~frank/sobweb1.pdf}


\bibitem{Frank08}
R.L. Frank, R. Seiringer,
Non-linear ground state representations and sharp Hardy inequalities.
\emph{J. Funct. Anal.} \textbf{255} (2008), no. 12, 3407--3430.


\bibitem{Frohlich86}
J. Fr\"{o}hlich, E.H. Lieb, M. Loss,
Stability of Coulomb systems with magnetic fields. I. The one-electron atom.
\emph{Comm. Math. Phys.} \textbf{104} (1986), no. 2, 251--270.


\bibitem{Lam20}
N. Lam, G. Lu, L. Zhang,
Geometric Hardy's inequalities with general distance functions.
\emph{J. Funct. Anal.} \textbf{279} (2020), no. 8, 108673, 35 pp.


\bibitem{Lin86}
C.-S. Lin,
Interpolation inequalities with weights.
\emph{Comm. Partial Differential Equations} \textbf{11} (1986), no. 14, 1515--1538.


\bibitem{Nguyen15}
V.H. Nguyen,
Sharp weighted Sobolev and Gagliardo-Nirenberg inequalities on half-spaces via mass transport and consequences.
\emph{Proc. Lond. Math. Soc. (3)} \textbf{111} (2015), no. 1, 127--148.


\bibitem{Talenti76}
G. Talenti,
Best constant in Sobolev inequality.
\emph{Ann. Mat. Pura Appl. (4)} \textbf{110} (1976), 353--372.


\bibitem{Tertikas07}
A. Tertikas, N.B. Zographopoulos,
Best constants in the Hardy-Rellich inequalities and related improvements.
\emph{Adv. Math.} \textbf{209} (2007), no. 2, 407--459.


\bibitem{Vazquez00}
J.L. Vazquez, E. Zuazua,
The Hardy inequality and the asymptotic behaviour of the heat equation with an inverse-square potential.
\emph{J. Funct. Anal.} \textbf{173} (2000), no. 1, 103--153.


\bibitem{Xia07}
C. Xia,
The Caffarelli-Kohn-Nirenberg inequalities on complete manifolds.
\emph{Math. Res. Lett.} \textbf{14} (2007), no. 5, 875--885.


\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\date{}
\makeatletter
\@addtoreset{equation}{section}
\makeatother

\usepackage[colorlinks,
            linkcolor=blue,
            anchorcolor=blue,
            citecolor=blue]{hyperref}
\usepackage{hyperref}
\hypersetup{colorlinks=true,
            linkcolor=blue,
            anchorcolor=blue,
            citecolor=blue}
\usepackage{hyperref,amsmath,amssymb,amscd}
\usepackage{amssymb,amsmath,enumerate}
\usepackage{indentfirst}
\usepackage{url}
\usepackage[capitalise]{cleveref}
\usepackage{amsthm}
\usepackage{amsmath}
\usepackage{diagbox}

\AtBeginDocument{\hypersetup{pdfborder={0 0 0.01}}}
\newtheorem{theorema}{Theorem}[section]

\newtheorem{thm}{Theorem}[section]

\newtheorem{lem}[thm]{Lemma}

\theoremstyle{definition}

\newtheorem{rem}[thm]{\bf Remark}
\allowdisplaybreaks[4]

\usepackage{enumitem}
\usepackage{extarrows}
\usepackage{amsthm}
\usepackage{setspace}
\hypersetup{urlcolor=blue, citecolor=red}
\allowdisplaybreaks[4]
\usepackage{cite}

\begin{document}

\title{\bf
On the extremal functions of second order uncertainty principles: symmetry and symmetry breaking\footnote{This work was supported by National Natural Science Foundation of China (No. 12371120).}}
\author{{Xiao-Ping Chen,\ \ Chun-Lei Tang\footnote{Corresponding author.\newline
\indent\,\,\, \emph{E-mail address:}   xpchen\_maths@163.com (X.-P. Chen); tangcl@swu.edu.cn (C.-L. Tang).}}\\
{\small \emph{School of Mathematics and Statistics, Southwest University,  Chongqing {\rm400715},}}\\
{\small \emph{People's Republic of China}}}
\maketitle
\baselineskip 17pt

\noindent {\bf Abstract}:
\ This paper focus on the symmetry and symmetry breaking about the second order Hydrogen Uncertainty Principle. \emph{Firstly}, by choosing a suitable test function, we give a negative answer to the conjecture presented by Cazacu, Flynn and Lam in [\emph{J. Funct. Anal.} \textbf{283} (2022), Paper No. 109659, 37 pp] for $N\in\{2,3\}$, and emphasizing the symmetry breaking phenomenon. \emph{Secondly}, we obtain a family of sharp weighted second order Hydrogen Uncertainty Principle, and prove the extremal functions are radial, which extends the work of Duong and Nguyen [The sharp second order Caffareli-Kohn-Nirenberg inequality and stability estimates for the sharp second order uncertainty principle,  arXiv:2102.01425].

\vspace{0.25em}

\noindent\textbf{Keywords}:\ Caffarelli-Kohn-Nirenberg inequalities; second order uncertainty principles; sharp constants; extremal functions

\vspace{0.25em}

\noindent\textbf{MSC}:
26D10 (primary); 46E35 (secondary)

\section{Introduction and main results}

\noindent This paper aims to study second order Hydrogen Uncertainty Principles (HyUP for short), which can be stated more precisely as follows.
\begin{enumerate}
[itemsep=0pt,
topsep=2pt,
parsep=2pt]

\item[($a$)]
Firstly, we consider the following second order HyUP,
\begin{equation*}
\int_{\mathbb{R}^N}|\Delta u|^2\mathrm{d}x
\int_{\mathbb{R}^N}\left|\nabla u\right|^2\mathrm{d}x
\ge C(N)
\left(\int_{\mathbb{R}^N}
\frac{\left|\nabla u\right|^2}{|x|}
\mathrm{d}x\right)^2,\quad \forall u\in C_c^\infty(\mathbb{R}^N),
\end{equation*}
which has been proved by Cazacu, Flynn and Lam in \cite[Theorem 2.3]{Cazacu22} for higher dimensions $N\geq 5$ with $C(N)=\frac{(N+1)^2}{4}$. Moreover, the authors in \cite[Remark 2.4]{Cazacu22} conjectured that
$C(N)=\frac{(N+1)^2}{4}$ for $2\leq N\leq 4$ . In Theorem \ref{thmnsup} below, by choosing a suitable test function, we will give a negative answer to this conjecture, i.e., $C(N)<\frac{(N+1)^2}{4}$ for $N\in\{2,3\}$, and emphasize the symmetry breaking phenomenon.

\item[($b$)]
Secondly, by using spherical harmonics decomposition, we study the following sharp weighted second order HyUP:
\begin{equation}\label{8.16-4}
\int_{\mathbb{R}^N}\frac{|\Delta u|^2}{|x|^{2\alpha}}\mathrm{d}x
\int_{\mathbb{R}^N}\left|\nabla u\right|^2\mathrm{d}x
\ge C(N,\alpha)
\left(\int_{\mathbb{R}^N}
\frac{\left|\nabla u\right|^2}{|x|^{\alpha+1}}
\mathrm{d}x\right)^2,\quad \forall u\in
C_c^\infty(\mathbb{R}^N),
\end{equation}
where
\begin{equation*}
C(N,\alpha)=
\begin{cases}
\frac{\alpha^2}{4},
\ \ &
\mbox{if}
\
N=1
\
\mbox{and}
\
-1<\alpha\le-\frac{1}{2},\\[1.5mm]
\frac{(3\alpha+2)^2}{4},
\ \ &
\mbox{if}
\
N=1
\
\mbox{and}
\
\alpha>-\frac{1}{2},\\[1.5mm]
\frac{(N+3\alpha+1)^2}{4},
\ \ &
\mbox{if}
\
N\ge2,
\
\alpha>-1
\
\mbox{and}
\
N\ge5\alpha+5;
\end{cases}
\end{equation*}
we also obtain the sufficient and necessary condition such that the equality in \eqref{8.16-4} holds, see Theorem \ref{thmwsupsc} below. These results improve and generalize that of \cite{Cazacu22}.

\end{enumerate}

The rest of this section is organized as follows. In Section \ref{sect-1.1}, we recall some significant results of the first order and second order HyUP (or some related inequalities). Our main results and related comments are described in Section \ref{sect-1.2}.

\subsection{Overview and motivation}\label{sect-1.1}

\noindent We begin with the first order HyUP that has important applications in mathematics and  physics (for example, which can be used to study the stability of a hydrogenic atom in a magnetic field, see e.g., \cite{Frohlich86}): for $N\ge2$,
\begin{equation}\label{8.11-8}
\int_{\mathbb{R}^N}
|\nabla u|^2\mathrm{d}x
\int_{\mathbb{R}^N}
\left|u\right|^2\mathrm{d}x
\ge\frac{(N-1)^2}{4}
\left(\int_{\mathbb{R}^N}
\frac{\left|u\right|^2}{|x|}
\mathrm{d}x\right)^2,
\end{equation}
where the constant $\frac{(N-1)^2}{4}$ in \eqref{8.11-8} is sharp and can be attained by $u(x)=ae^{-b|x|}$ with $a\in\mathbb{R}$ and $b>0$, see \cite[Example 1.6]{Frank11} for details.

The HyUP \eqref{8.11-8} is a special case of the first order interpolation inequalities which is known as   Caffarelli-Kohn-Nirenberg (CKN for short) inequalities (see \cite{Caffarelli84}). A large number of scholars have paid attention to CKN inequalities in the past several decades due to their applications in many area of mathematics and physics, we refer the readers to see \cite{Catrina09,Catrina01,Cazacu21,
Cazacu24-2,Costa08,Dong18,
Xia07,Do23,Duy22,Frank08,
Lam20,Vazquez00,DelPino03,
Nguyen15,Talenti76,Brezis18,
Cordero-Erausquin04,DelPino02} for first order CKN inequalities and the references therein.

Compared with the first order CKN inequalities, there exist few results about the higher order CKN inequalities except \cite{Cazacu22,Cazacu23,
Duong21,Duong25,Lin86}. We first introduce some basic notations in order to state the previous results about higher order CKN inequalities. For $\alpha,\beta$ satisfying $N-2\alpha>0$ and $N-2\beta>0$, let $H^2_{\alpha,\beta}
(\mathbb{R}^N)$ be the second order Sobolev space which is the closure of $C_c^\infty(\mathbb{R}^N)$ with respect to the norm
\begin{equation*}
\|u\|_{H^2_{\alpha,\beta}
(\mathbb{R}^N)}
:=\left(\int_{\mathbb{R}^N}
\frac{|\Delta u|^2}
{|x|^{2\alpha}}\mathrm{d}x
+\int_{\mathbb{R}^N}
\frac{|\nabla u|^2}
{|x|^{2\beta}}\mathrm{d}x
\right)^{\frac{1}{2}}.
\end{equation*}

In the recent paper \cite{Cazacu22}, Cazacu, Flynn and Lam considered the following  second order HyUP in higher dimensions $N\ge5$, and proved the following results.

\begin{theorema}
[{\!\rm{\!\cite[Theorem 2.3]{Cazacu22}}}]
\label{thm-c}
Let $N\ge5$. Then the following inequality holds
\begin{equation}\label{sUP}
\int_{\mathbb{R}^N}
|\Delta u|^2
\mathrm{d}x
\int_{\mathbb{R}^N}
|\nabla u|^2\mathrm{d}x
\geq C(N)
\left(\int_{\mathbb{R}^N}
\frac{|\nabla u|^2}{|x|}
\mathrm{d}x\right)^2,
\end{equation}
for any $u\in
H^2_{0,0}(\mathbb{R}^N)$, where $C(N)=\frac{(N+1)^2}{4}$ is the sharp constant of \eqref{sUP} and can be attained by $u(x)=a(1+b|x|)\exp(-b|x|)$ with   $a\in\mathbb{R}$ and $b>0$.
\end{theorema}

Note that their main idea is to apply spherical harmonics decomposition, but their proof fails for lower dimensions $2\le N\le4$. Moreover, let us denote $\mathcal{R}_1:
=\frac{x}{|x|}\cdot\nabla$ and $\mathcal{R}_2:
=\mathcal{R}_1^2
+\frac{N-1}{r}\mathcal{R}_1$, where $\mathcal{R}_2:
=\mathcal{R}_1\circ\mathcal{R}_1$, which represent the radial derivative and the radial Laplacian, respectively. Then, in \cite[Theorem 2.4]{Cazacu22}, the authors also proved that, for all  $N\geq 2$,
\begin{equation*}
\int_{\mathbb{R}^N}
|\mathcal{R}_2 u|^2\mathrm{d}x
\int_{\mathbb{R}^N}
\left|\mathcal{R}_1 u\right|^2 \mathrm{d}x
\ge \frac{(N+1)^2}{4}
\left(\int_{\mathbb{R}^N}
\frac{\left|\mathcal{R}_1 u\right|^2}{|x|}
\mathrm{d}x\right)^2,
\end{equation*}
and the equality holds by $u(x)=a(1+b|x|)e^{-b|x|}$ with  $a\in\mathbb{R}$ and $b>0$. That is to say, for any radial function $u\in
H^2_{0,0}(\mathbb{R}^N)$, $C(N)=\frac{(N+1)^2}{4}$ for $2\le N\le4$. Therefore, for all $u\in
H^2_{0,0}(\mathbb{R}^N)$ (without radiality assumption), $C(N)\le\frac{(N+1)^2}{4}$ for $2\le N\le4$. \emph{In} \cite[Remark 2.4]{Cazacu22}, \emph{the authors conjectured that the sharp constant in \eqref{sUP} satisfies $C(N)
=\frac{(N+1)^2}{4}$ for $2\leq N\leq 4$}. We will give a negative answer to this conjecture for $N\in\{2,3\}$ in Theorem \ref{thmnsup} below.

It is worthy to mention that the authors in \cite{Cazacu22} also studied the so-called Heisenberg Uncertainty Principle, and obtained the following results.

\begin{theorema}
[{\!\rm{\!\cite[Theorem 2.1]{Cazacu22}}}]
\label{thm-d}
Let $N\ge1$. Then for any $u\in
H^2_{0,-1}(\mathbb{R}^N)$,
\begin{equation}\label{sUP-2}
\int_{\mathbb{R}^N}
|\Delta u|^2
\mathrm{d}x
\int_{\mathbb{R}^N}
|x|^2|\nabla u|^2\mathrm{d}x
\geq \frac{(N+2)^2}{4}
\left(\int_{\mathbb{R}^N}
|\nabla u|^2\mathrm{d}x\right)^2,
\end{equation}
where $\frac{(N+2)^2}{4}$ is the sharp constant of \eqref{sUP-2} and is attained by $u(x)=a\exp(-b|x|^2)$ with  $a\in\mathbb{R}$ and $b>0$.
\end{theorema}

Subsequently, Cazacu, Flynn and Lam in \cite{Cazacu23} improved \eqref{sUP} and \eqref{sUP-2} to the following  weighted version.

\begin{theorema}
[{\!\rm{\!\cite[Corollary 2.4]{Cazacu23}}}]
\label{thm-b}
Let $N\ge5$. The following inequality
\begin{equation}\label{8.11-6}
\int_{\mathbb{R}^N}
|\Delta u|^2
\mathrm{d}x
\int_{\mathbb{R}^N}
\frac{|\nabla u|^2}
{|x|^{2\beta}}\mathrm{d}x
\geq C(N,0,\beta)
\left(\int_{\mathbb{R}^N}
\frac{|\nabla u|^2}
{|x|^{\beta+1}}
\mathrm{d}x\right)^2
\end{equation}
holds for any $u\in H^2_{0,\beta}(\mathbb{R}^N)$. Furthermore,
\begin{enumerate}
[itemsep=0pt,
topsep=2pt,
parsep=2pt]

\item[(1)]
if $\beta<1$, then $C(N,0,\beta)
=\frac{(N-\beta+1)^2}{4}$ is the sharp constant of \eqref{8.11-6} and can be attained by the functions $u$ satisfying $\nabla u=a\exp\left(-\frac{b}{1-\beta}
|x|^{1-\beta}\right)x$ with  $a\in\mathbb{R}$ and $b>0$;

\item[(2)]
if $\beta>1$, then $C(N,0,\beta)
=\frac{(N+\beta-1)^2}{4}$ is the sharp constant of \eqref{8.11-6} and can be attained by the functions $u$ satisfying $\nabla u=a|x|^{-N} \exp\left(-\frac{b}{1-\beta}
|x|^{1-\beta}\right)x$ with $a\in\mathbb{R}$ and $b<0$.

\end{enumerate}
\end{theorema}

In particular, when $\beta=1$ in \eqref{8.11-6}, it becomes the well-known sharp Hardy-Rellich inequality, see \cite{Tertikas07,Beckner08,Cazacu20} for details.

Along the same line of thought, the authors in \cite{Cazacu23} also obtained the following sharp weighted second order Heisenberg Uncertainty Principle.

\begin{theorema}
[{\!\rm{\!\cite[Corollary 2.5]{Cazacu23}}}]
\label{thm-a}
Let $N\ge2$ and $\alpha\in\mathbb{R}$. There holds
\begin{equation}\label{gwsup-1}
\int_{\mathbb{R}^N}
\!\frac{|\Delta u|^2}
{|x|^{2\alpha}}
\mathrm{d}x
\int_{\mathbb{R}^N}
|x|^{2\alpha+2}|\nabla u|^2\mathrm{d}x
\geq \frac{(N+4\alpha+2)^2}{4}
\left(\int_{\mathbb{R}^N}
|\nabla u|^2
\mathrm{d}x\right)^2,
\end{equation}
for any $u\in
H^2_{\alpha,-\alpha-1}
(\mathbb{R}^N)$. Furthermore, if either $\alpha+1>0$ or $N+4\alpha+2>0$, then $\frac{(N+4\alpha+2)^2}{4}$ is the sharp constant of \eqref{gwsup-1} and can be attained by $u(x)=a
\exp\left[-b|x|^{2(\alpha+1)}\right]$ with $a\in\mathbb{R}$ and $b>0$.
\end{theorema}

Recently, Duong and Nguyen in \cite{Duong21} established a family of more general second order CKN inequalities: let $N\geq 1$ and $\alpha,\beta$ be such that
\begin{equation*}
N-2\alpha>0,
\
N-2\beta>0,
\
N-\alpha-\beta-1>0,
\end{equation*}
then
\begin{equation}\label{gwsup}
\int_{\mathbb{R}^N}
\frac{|\Delta u|^2}
{|x|^{2\alpha}}
\mathrm{d}x
\int_{\mathbb{R}^N}
\frac{|\nabla u|^2}
{|x|^{2\beta}}\mathrm{d}x
\geq C(N,\alpha,\beta)
\left(\int_{\mathbb{R}^N}
\frac{|\nabla u|^2}
{|x|^{\alpha+\beta+1}}
\mathrm{d}x\right)^2,
\end{equation}
for any $u\in H^2_{\alpha,\beta}
(\mathbb{R}^N)$. Furthermore, if $\alpha-\beta+1>0$ and $N+2\alpha>0$, the constant $C(N,\alpha,\beta)$ in \eqref{gwsup} is sharp and  satisfies
\begin{equation*}
C(N,\alpha,\beta)
\geq\inf_{k\in\mathbb{N}\cup\{0\}}
\dfrac{1+\min\left\{0,
\frac{8\beta k}
{\left(N+2k-2\beta-2\right)^2}
\right\}}
{\left[1+\max\left\{0,
\frac{4(\alpha+\beta+1)k}
{\left(N+2k-\alpha-\beta-3\right)^2
}\right\}\right]^2}
\left(\frac{N+2k+3\alpha-\beta+1}
{2}\right)^2.
\end{equation*}
However, the authors did not give the explicit form of $C(N,\alpha,\beta)$ and its extremal functions.

Note that \eqref{gwsup} contains \eqref{sUP}, \eqref{sUP-2}, \eqref{8.11-6} and \eqref{gwsup-1}. For the sake of comparison, we present the previous results of \eqref{gwsup} about sharp constants and extremal functions in the table below:

\vspace*{0.25em}

\begin{center}
\begin{tabular}
{l|c|c|c}
  \hline
  \diagbox
  {$\alpha,\beta$}
  {exist or not}
  {$N$}
  &
  $N=2$
  &
  $N=3,4$
  &
  $N\ge5$
  \\
  \hline
  $\alpha=0$, $\beta=0$
  &

  &

  &
  \cite[Theorem 2.3]{Cazacu22}
  \\
  \hline
  $\alpha=0$, $\beta=1$
  &

  &
  \cite{Beckner08},
  \cite[Theorem 1.1]{Cazacu20}
  &
  \cite[(1.8)]{Tertikas07},
  \cite[Theorem 1.1]{Cazacu20}
  \\
  \hline
  $\alpha=0$, $\beta=-1$
  &
  \cite[Theorem 2.1]{Cazacu22}
  &
  \cite[Theorem 2.1]{Cazacu22}
  &
  \cite[Theorem 2.1]{Cazacu22}
  \\
  \hline
  $\alpha=0$, $\beta\in\mathbb{R}$
  &

  &

  &
  \cite[Corollary 2.4]{Cazacu23}
  \\
  \hline
  $\alpha\in\mathbb{R}$,
  $\beta=-\alpha-1$
  &
  \cite[Corollary 2.5]{Cazacu23}
  &
  \cite[Corollary 2.5]{Cazacu23}
  &
  \cite[Corollary 2.5]{Cazacu23}
  \\
  \hline
\end{tabular}
\end{center}

\vspace*{0.5em}

Inspired by the papers mentioned above, it is natural to ask a question that \emph{when $\alpha\in\mathbb{R}$ and $\beta=0$ in \eqref{gwsup}, what is the explicit form of $C(N,\alpha,0)$ and its extremal functions}. We will give a positive answer of it in Theorem \ref{thmwsupsc} below.

\subsection{Main results and some comments}\label{sect-1.2}

\noindent The first result of this paper reads as follows.

\begin{thm}\label{thmnsup}
For $N\in\{2,3\}$, the sharp constant in \eqref{sUP} satisfies
\begin{equation}\label{nsup}
C(N)<\frac{(N+1)^2}{4},
\end{equation}
more precisely,
\begin{equation*}
\frac{(N-1)^4(N+3)^2}
{4(N^2-2N+5)^2}
\leq C(N)\leq \frac{N(N+4)(N^2-1)^2}
{4(N^2-N+4)^2}
<\frac{(N+1)^2}{4}.
\end{equation*}
\end{thm}

\begin{rem}
Here we present some comments about Theorem \ref{thmnsup}.
\begin{enumerate}
[itemsep=0pt,
topsep=2pt,
parsep=2pt]

\item[(1)]
The key point of proving Theorem \ref{thmnsup} is to choose a suitable test function $u_*\in C_c^\infty(\mathbb{R}^N)$ such that
\begin{equation*}
\frac{\int_{\mathbb{R}^N}
|\Delta u_*|^2\mathrm{d}x
\int_{\mathbb{R}^N}
\left|\nabla u_*\right|^2 \mathrm{d}x}
{\left(\int_{\mathbb{R}^N}
\frac{
\left|\nabla u_*\right|^2}{|x|}
\mathrm{d}x\right)^2}
<\frac{(N+1)^2}{4},
\quad
\mbox{for}
\
N\in\{2,3\}.
\end{equation*}
When $N=4$, it seems not easy to choose such suitable test function, and thus we can not obtain \eqref{nsup}. We conjecture that $\eqref{nsup}$ also holds for $N=4$.

\item[(2)]
From Theorem \ref{thmnsup}, we know that when $N\in\{2,3\}$, if the extremal function of \eqref{sUP} exists, it must be non-radial, this can be called symmetry breaking phenomenon.

\end{enumerate}
\end{rem}

Our second result is given in the following theorem.

\begin{thm}\label{thmwsupsc}
Assume that $N\ge1$ and $\alpha>-1$. For any $u\in H^2_{\alpha,0}(\mathbb{R}^N)$,  there hold:
\begin{enumerate}
[itemsep=0pt,
topsep=2pt,
parsep=2pt]

\item[(1)] Case 1: $N=1$.

\begin{enumerate}
[itemsep=0pt,
topsep=0pt,
parsep=0pt]

\item[(1a)]
if $-1<\alpha\le-\frac{1}{2}$, then
\begin{equation}\label{8.15-1}
\int_{\mathbb{R}}
\frac{|\Delta u|^2}
{|x|^{2\alpha}}
\mathrm{d}x
\int_{\mathbb{R}}\left|\nabla u\right|^2\mathrm{d}x
\ge\frac{\alpha^2}{4}
\left(\int_{\mathbb{R}}
\frac{\left|\nabla u\right|^2}
{|x|^{\alpha+1}}
\mathrm{d}x\right)^2,
\end{equation}
where the constant $\frac{\alpha^2}{4}$ is sharp and is attained if and only if $u(x)=a\int^{\infty}_{|x|}
\exp\left(-br^{\alpha+1}\right)
\mathrm{d}r$ with $a\in \mathbb{R}$ and $b>0$;

\item[(1b)]
if $\alpha>-\frac{1}{2}$ or $\alpha<-1$, then
\begin{equation}\label{8.15-2}
\int_{\mathbb{R}}
\frac{|\Delta u|^2}
{|x|^{2\alpha}}
\mathrm{d}x
\int_{\mathbb{R}}\left|\nabla u\right|^2\mathrm{d}x
\ge\frac{(3\alpha+2)^2}{4}
\left(\int_{\mathbb{R}}
\frac{\left|\nabla u\right|^2}
{|x|^{\alpha+1}}
\mathrm{d}x\right)^2,
\end{equation}
where the constant $\frac{(3\alpha+2)^2}{4}$ is sharp and is attained if and only if  $u(x)=a(1+b|x|^{\alpha+1})
\exp\left(-b |x|^{\alpha+1}\right)$ with $a\in\mathbb{R}$ and $b>0$.

\end{enumerate}

\item[(2)]
Case 2: $N\ge2$. Under the additional assumptions $N\geq 5\alpha+5$, then
\begin{equation}\label{wsUP}
\int_{\mathbb{R}^N}
\frac{|\Delta u|^2}
{|x|^{2\alpha}}
\mathrm{d}x
\int_{\mathbb{R}^N}\left|\nabla u\right|^2\mathrm{d}x
\ge\frac{(N+3\alpha+1)^2}{4}
\left(\int_{\mathbb{R}^N}
\frac{\left|\nabla u\right|^2}
{|x|^{\alpha+1}}
\mathrm{d}x\right)^2,
\end{equation}
where the constant $\frac{(N+3\alpha+1)^2}{4}$ is sharp and is attained if and only if  $u(x)=a(1+b|x|^{\alpha+1})
\newline
\exp\left(-b |x|^{\alpha+1}\right)$ with $a\in \mathbb{R}$ and $b>0$.

\end{enumerate}
\end{thm}

\begin{rem}
Some discussions about Theorem \ref{thmwsupsc} are presented below.
\begin{enumerate}
[itemsep=0pt,
topsep=2pt,
parsep=2pt]

\item[(1)]
When $\alpha=0$ in Theorem \ref{thmwsupsc}-$(1b)$, the inequality \eqref{8.15-2} reduces into
\begin{equation*}
\int_{\mathbb{R}}
|\Delta u|^2
\mathrm{d}x
\int_{\mathbb{R}}\left|\nabla u\right|^2\mathrm{d}x
\ge
\left(\int_{\mathbb{R}}
\frac{\left|\nabla u\right|^2}
{|x|}
\mathrm{d}x\right)^2,
\end{equation*}
and it is attained by  $u(x)=a(1+b|x|)
e^{-b |x|}$ with $a\in \mathbb{R}$ and $b>0$, which indicates that the result of \cite[Theorem 2.3]{Cazacu22} (or see Theorem \ref{thm-c} above) also holds for $N=1$. Therefore, combining with Theorem \ref{thmnsup}, the conjecture in \cite[Remark 2.4]{Cazacu22} only remains unsolved when $N=4$.

\item[(2)]
Compared with \cite[Theorem 2.3]{Cazacu22} (or see Theorem \ref{thm-c} above), we obtain a weighted version of the second order HyUP \eqref{sUP} in Theorem \ref{thmwsupsc}-(2), more precisely, if $\alpha=0$ in Theorem \ref{thmwsupsc}-(2), it reduces to \cite[Theorem 2.3]{Cazacu22}.

\item[(3)]
We also obtain the sufficient and necessary conditions such that the equality in \eqref{8.15-1}$-$\eqref{wsUP} hold.

\end{enumerate}
\end{rem}

The rest of this paper is organized as follows. In Section \ref{sect-2}, we aim to analyze the symmetry breaking phenomenon for the second order HyUP, and prove Theorem \ref{thmnsup}. In Section \ref{sect-3}, we establish a family of weighted second order HyUP, and prove Theorem \ref{thmwsupsc}.

Throughout this paper, we will frequently use the following notations.
\begin{itemize}
\setlength{\itemsep}{0pt}
\setlength{\parsep}{0pt}
\setlength{\parskip}{2pt}

\item
$\mathbb{N}:=\{1,2,\cdots\}$ denotes the set which includes all natural numbers.

\item
$\mathbb{S}^{N-1}$ denotes the $(N\!-\!1)$-dimensional sphere with respect to the Hausdorff measure in $\mathbb{R}^N$.

\item $o_\varepsilon(1)$ denotes a quantity which tends to zero as $\varepsilon\to 0^+$.

\end{itemize}

\section{Symmetry breaking phenomenon for the second order HyUP: proof of Theorem \ref{thmnsup}}\label{sect-2}

\noindent In this section, we will show a symmetry breaking phenomenon for the second order HyUP \eqref{sUP}, that is, $C(N)<\frac{(N+1)^2}{4}$ for $N\in\{2,3\}$. Notice that the symmetry result heavily depends on the following conclusion obtained by Cazacu, Flynn and Lam  \cite[Lemma 3.4]{Cazacu22}: for any $N\geq 5$, \begin{equation*}
\inf_{k\in\mathbb{N}\cup\{0\}}
\frac{1}{\left[1+\frac{4k}
{(N+2k-3)^2}\right]^2}
\frac{(N+2k+1)^2}{4}
=\frac{(N+1)^2}{4}.
\end{equation*}
For $2\le N\le4$, we present the following calculation.

\begin{lem}\label{lemtr}
For all $2\leq N\leq 4$, it holds that
\begin{equation*}
\inf_{k\in\mathbb{N}\cup\{0\}}
\frac{1}{\left[1+\frac{4k}
{(N+2k-3)^2}\right]^2}
\frac{(N+2k+1)^2}{4}
=\frac{(N-1)^4(N+3)^2}
{4(N^2-2N+5)^2}.
\end{equation*}
\end{lem}

\begin{proof}[\rm\textbf{Proof}]
Note that
\[
\frac{1}{\left[1+\frac{4k}
{(N+2k-3)^2}\right]^2}
\frac{(N+2k+1)^2}{4}
=\frac{(N+2k-3)^4}
{[(N+2k-3)^2+4k]^2}
\frac{(N+2k+1)^2}{4}
:=\mathcal{J}(N,k).
\]

$\bullet$ \emph{Step 1: For $2\le N\le 4$, we will prove that}
\[
\inf_{k\in\mathbb{N}}
\mathcal{J}(N,k)
=\mathcal{J}(N,1).
\]

For $2\le N\le4$, let us define the following function on the real axis $[2,\infty)$:
\begin{align*}
f(x)
:=\frac{(N+x-3)^4}{[(N+x-3)^2+2x]^2}
\frac{(N+x+1)^2}{4}
=g(x)h(x),
\end{align*}
where
\begin{align*}
g(x):&=\frac{(N+x+1)^2}
{4[x^2+2(N-2)x+(N-3)^2
]^{\frac{1}{4}}}>0,
\\
h(x):&=\frac{(N+x-3)^4}
{[x^2+2(N-2)x+(N-3)^2
]^{\frac{7}{4}}}>0.
\end{align*}
It is easy to check that
\begin{align*}
g'(x)&
=\frac{(N+x+1)
\left[3x^2+3(2N-5)x+4(N-3)^2
-(N^2-N-2)\right]}
{8[x^2+2(N-2)x+(N-3)^2
]^{\frac{5}{4}}},
\\
h'(x)&=\frac{(N+x-3)^3
\left[x^2+(2N+3)x+(N^2-13N+30)
\right]}
{2[x^2+2(N-2)x+(N-3)^2
]^{\frac{11}{4}}}.
\end{align*}

Under the condition $2\le N\le 4$, it is easy to get that  $-\frac{2N-5}{2}
\le\frac{1}{2}<2$, and then for all $x\ge2$,
\begin{align*}
&3x^2+3(2N-5)x+4(N-3)^2-(N^2-N-2)
\\&\quad\ge
12+6(2N-5)+4(N-3)^2-(N^2-N-2)
\\&\quad=
3N^2-11N+20>0.
\end{align*}
This leads to, $g'(x)>0$ for all $x\geq 2$.

On the other hand, using $2\le N\le4$ again,  $-\frac{2N+3}{2}<0<2$ obviously, and then for all $x\ge2$,
\begin{align*}
&x^2+(2N+3)x+(N^2-13N+30)
\\&\quad\ge
4+2(2N+3)+(N^2-13N+30)
\\&\quad=N^2-9N+40>0.
\end{align*}
This implies that $h'(x)>0$ for all $x\geq 2$.

To sum up,
\[
f'(x)=g'(x)h(x)+g(x)h'(x)>0
\]
for all $x\geq 2$, this means that $f(x)$ is increasing for all $x\geq 2$. Thus, $f(x)\ge f(2)$ for all $x\ge2$, then for all $k\in\mathbb{N}$,
\begin{equation*}
\mathcal{J}(N,k)
=f(2k)\ge f(2)
=\mathcal{J}(N,1),
\end{equation*}
namely,
\begin{equation*}
\inf_{k\in \mathbb{N}}
\mathcal{J}(N,k)
=\mathcal{J}(N,1).
\end{equation*}
The proof of \emph{Step 1} finishes.

From \emph{Step 1}, we see that
\begin{align}\label{8.16-1}
\inf_{k\in\mathbb{N}\cup\{0\}}
\mathcal{J}(N,k)
&=\min\{\mathcal{J}(N,0),
\mathcal{J}(N,1)\}.
\end{align}

$\bullet$ \emph{Step 2: For $2\le N\le 4$, we will verify that}
\begin{equation}\label{8.16-2}
\mathcal{J}(N,1)
<\mathcal{J}(N,0).
\end{equation}

Notice that
\begin{eqnarray*}
\mathcal{J}(N,1)
=\frac{(N-1)^4(N+3)^2}
{4(N^2-2N+5)^2}
=\left\{\arraycolsep=1.5pt
\begin{array}{ll}
\frac{1}{4},
\quad &\mbox{if}\  N=2,\\[2mm]
\frac{9}{4},
\quad &\mbox{if}\  N=3,\\[2mm]
\frac{3969}{676}\approx 5.87, \quad &\mbox{if}\  N=4,
\end{array}
\right.
\end{eqnarray*}
and
\begin{eqnarray*}
\mathcal{J}(N,0)
=\frac{(N+1)^2}{4}
=\left\{\arraycolsep=1.5pt
\begin{array}{ll}
\frac{9}{4},
\quad &\mbox{if}\  N=2,\\[2mm]
4,
\quad &\mbox{if}\  N=3,\\[2mm]
\frac{25}{4}, \quad &\mbox{if}\  N=4,
\end{array}
\right.
\end{eqnarray*}
then \eqref{8.16-2} holds. This ends the proof of \emph{Step 2}.

$\bullet$ \emph{Step 3: Conclusion.}

From \eqref{8.16-1} and \eqref{8.16-2},
\[
\inf_{k\in\mathbb{N}\cup\{0\}}
\mathcal{J}(N,k)
=\mathcal{J}(N,1)
=\frac{(N-1)^4(N+3)^2}
{4(N^2-2N+5)^2},
\]
as our desired estimate. The proof is completed.
\end{proof}

Now, it remains to prove Theorem \ref{thmnsup} to complete this section. The main idea is to adapt the well-known technique of decomposing a function $u$ into spherical harmonicas, which is a useful method (see \cite{Cazacu22,Duong21,Cazacu20,
Tertikas07,Vazquez00} and the references therein). For $N\ge2$, we apply the coordinate transformation $x\in\mathbb{R}^N
\mapsto(r,\sigma)
\in(0,\infty)\times
\mathbb{S}^{N-1}$, and then for each $u\in C^\infty_c(\mathbb{R}^N)$,
\begin{equation}\label{8.13-1}
u(x)=u(r\sigma)
=\sum^\infty_{k=0}
u_k(r)\phi_k(\sigma),
\end{equation}
where $\phi_k$ are the orthonormal eigenfunctions of the Laplace-Beltrami operator on $\mathbb{S}^{N-1}$ with the corresponding eigenvalue $c_k=k(N+k-2)$, namely, $-\Delta_{\mathbb{S}^{N-1}}\phi_k
=c_k\phi_k$ with $k\in\mathbb{N}\cup\{0\}$. The Fourier coefficients $\{u_k\}_k$ belong to
$C_c^\infty([0,\infty))$ and satisfy $u_{k}(r)=O(r^k)$ as $r\to 0$, which allows us to consider the change of variables
\begin{equation*}
u_{k}(r)=r^kv_k(r),
\end{equation*}
where $v_k\in C^\infty_c([0,\infty))$ (see \cite[page 418]{Tertikas07} for more details).

Now, based on Lemma \ref{lemtr} and above properties, we are in a position to prove Theorem \ref{thmnsup}.

\begin{proof}
[\rm\textbf{Proof of Theorem \ref{thmnsup}}]

From the work of Cazacu, Flynn and  Lam \cite[Proof of Theorem 2.3]{Cazacu22}, we know that  the sharp constant of the second order HyUP \eqref{sUP} satisfies
\begin{equation*}
C(N)\geq \inf_{k\in\mathbb{N}\cup\{0\}}
\frac{1}{\left[1+\frac{4k}
{(N+2k-3)^2}\right]^2}
\frac{(N+2k+1)^2}{4},
\quad \mbox{for all}\ N\geq 2.
\end{equation*}
This together with Lemma \ref{lemtr} implies that, for $2\leq N\leq 4$,
\begin{equation*}
C(N)\geq
\frac{(N-1)^4(N+3)^2}
{4(N^2-2N+5)^2}.
\end{equation*}
Then, in order to prove the symmetry breaking phenomenon, that is, $C(N)<\frac{(N+1)^2}{4}$ for $N\in\{2,3\}$, it is enough to show that there exists a function $u_*\in C_c^\infty(\mathbb{R}^N)$ such that
\begin{equation}\label{sUPp}
\frac{\int_{\mathbb{R}^N}
|\Delta u_*|^2\mathrm{d}x
\int_{\mathbb{R}^N}
\left|\nabla u_*\right|^2 \mathrm{d}x}
{\left(
\int_{\mathbb{R}^N}
\frac{\left|\nabla u_*\right|^2}
{|x|}
\mathrm{d}x\right)^2}
<\frac{(N+1)^2}{4},
\quad
\mbox{for}
\
N\in\{2,3\}.
\end{equation}
Here, we choose a test function $u_*(x)=|x|e^{-|x|}\phi_1(\sigma)$, where $\phi_1$ is defined by \eqref{8.13-1} above. Then from \cite[Lemma 3.1]{Cazacu22} with  $v_1:=e^{-r}$, we have
\begin{align*}
\int_{\mathbb{R}^N}|\Delta u_*|^2\mathrm{d}x
&=\left|\mathbb{S}^{N-1}\right|
\left[\int_0^\infty|v_1''|^2 r^{N+1}\mathrm{d}r
+(N+1)\int_0^\infty|v_1'|^2 r^{N-1}\mathrm{d}r\right]
\\&=\left|\mathbb{S}^{N-1}\right|
\left[\int_0^\infty e^{-2r} r^{N+1}\mathrm{d}r
+(N+1)\int_0^\infty e^{-2r} r^{N-1}\mathrm{d}r\right]
\\&=\left|\mathbb{S}^{N-1}\right|
\left[\frac{\Gamma(N+2)}{2^{N+2}}
+(N+1)\frac{\Gamma(N)}{2^{N}}
\right],
\end{align*}
also
\begin{align*}
\int_{\mathbb{R}^N}
\left|\nabla u_*\right|^2 \mathrm{d}x
&=\left|\mathbb{S}^{N-1}\right|
\int_0^\infty|v_1'|^2 r^{N+1}\mathrm{d}r
\\&=\left|\mathbb{S}^{N-1}\right|
\int_0^\infty e^{-2r} r^{N+1}\mathrm{d}r
\\&=\left|\mathbb{S}^{N-1}\right|
\frac{\Gamma(N+2)}{2^{N+2}},
\end{align*}
and
\begin{align*}
\int_{\mathbb{R}^N}
\frac{\left|\nabla u_*\right|^2}{|x|}
\mathrm{d}x
&=\left|\mathbb{S}^{N-1}\right|
\left(\int_0^\infty|v_1'|^2 r^{N}\mathrm{d}r
+\int_0^\infty|v_1|^2 r^{N-2}\mathrm{d}r\right)
\\&=\left|\mathbb{S}^{N-1}\right|
\left(\int_0^\infty e^{-2r} r^{N}\mathrm{d}r
+\int_0^\infty e^{-2r} r^{N-2}\mathrm{d}r\right)
\\&=\left|\mathbb{S}^{N-1}\right|
\left[\frac{\Gamma(N+1)}{2^{N+1}}
+\frac{\Gamma(N-1)}{2^{N-1}}\right].
\end{align*}
Therefore,
\begin{align*}
\frac{\int_{\mathbb{R}^N}|\Delta u_*|^2\mathrm{d}x
\int_{\mathbb{R}^N}
\left|\nabla u_*\right|^2 \mathrm{d}x}{
\left(\int_{\mathbb{R}^N}
\frac{\left|\nabla u_*\right|^2}
{|x|}
\mathrm{d}x\right)^2}
&=\frac{
\left[\frac{\Gamma(N+2)}{2^{N+2}}
+(N+1)\frac{\Gamma(N)}{2^{N}}
\right]
\frac{\Gamma(N+2)}{2^{N+2}}}
{\left[\frac{\Gamma(N+1)}{2^{N+1}}
+\frac{\Gamma(N-1)}{2^{N-1}}
\right]^2}
\\&=\frac{N(N+4)(N^2-1)^2}
{4(N^2-N+4)^2},
\end{align*}
with the help of the property of the Gamma function: $\Gamma(t+1)=t\Gamma(t)$ for $t>0$. One can easily check that  $\frac{N(N+4)(N^2-1)^2}
{4(N^2-N+4)^2}
<\frac{(N+1)^2}{4}$ for $N\in\{2,3\}$, and then  \eqref{sUPp} holds. Thereby, the proof of Theorem \ref{thmnsup} is completed.
\end{proof}

\section{Symmetry phenomenon for the weighted second order HyUP: proof of Theorem \ref{thmwsupsc}}
\label{sect-3}

\noindent In this section, we analyze the sharp weighted second order HyUP, and prove Theorem \ref{thmwsupsc}. We first prove the case $N=1$ of Theorem \ref{thmwsupsc}.

\begin{proof}[\rm\textbf{Proof of Theorem \ref{thmwsupsc}: the case $\boldsymbol{N=1}$}]
We divide the proof into two cases.

$-$ \emph{Case $(1a)$: $-1<\alpha\le-\frac{1}{2}$}.

For all $u\in C_c^\infty(-\infty,\infty)$, \eqref{8.15-1} is equivalent to
\begin{align}\label{1Dc}
\int_{-\infty}^\infty
|u''|^2|x|^{-2\alpha}\mathrm{d}x
\int_{-\infty}^\infty
|u'|^2\mathrm{d}x
\ge\frac{\alpha^2}{4}
\left(\int_{-\infty}^\infty
|u'|^2|x|^{-\alpha-1}\mathrm{d}x
\right)^2.
\end{align}
Let us denote $w:=u'$, then \eqref{1Dc} becomes, for all $w\in C_c^\infty(-\infty,\infty)$,
\begin{align}\label{1Dcb}
\int_{-\infty}^\infty
|w'|^2|x|^{-2\alpha}\mathrm{d}x
\int_{-\infty}^\infty
|w|^2\mathrm{d}x
\ge\frac{\alpha^2}{4}
\left(\int_{-\infty}^\infty
|w|^2|x|^{-\alpha-1}\mathrm{d}x
\right)^2.
\end{align}
It was shown in
\cite[Lemma 3.4]{Cazacu24-2} (or
\cite[Corollary 1.1-(1)]{Do23}) that \eqref{1Dcb} holds, and the extremal functions are also  known. However, since our proof is different from those in \cite{Cazacu24-2,Do23}, let us give a simple proof here.

Under the assumption $-1<\alpha\leq -\frac{1}{2}$, using H\"{o}lder inequality and integration by parts,
\begin{align*}
\int_{-\infty}^\infty
|w'|^2|x|^{-2\alpha}\mathrm{d}x
\int_{-\infty}^\infty
|w|^2\mathrm{d}x
& \ge \left(\int_{-\infty}^\infty
w'w|x|^{-\alpha-1}x
\mathrm{d}x\right)^2
=\frac{\alpha^2}{4}
\left(\int_{-\infty}^\infty
|w|^2|x|^{-\alpha-1}
\mathrm{d}x\right)^2,
\end{align*}
and the first equality holds if and only if
\begin{align*}
|w'|^2|x|^{-2\alpha}
=C_1|w|^2
&\Rightarrow
w(x)=C_2
\exp\left(-c|x|^{\alpha+1}\right)
\\&\Rightarrow
u'(x)=C_2
\exp\left(-c|x|^{\alpha+1}\right),
\end{align*}
with the aid of $w=u'$, for some $c,C_1>0$ and $C_2\in\mathbb{R}$. Therefore, the equality in \eqref{1Dc} holds if and only if $u(x)=a\int^{\infty}_{|x|}
\exp\left(-b r^{\alpha+1}\right)
\mathrm{d}r$ with $a\in \mathbb{R}$ and $b>0$.

$-$ \emph{Case $(1b)$: $\alpha>-\frac{1}{2}$}.

For all $u\in C_c^\infty(-\infty,\infty)$, \eqref{8.15-2} is equivalent to
\begin{align}\label{1Dc2}
\int_{-\infty}^\infty
|u''|^2|x|^{-2\alpha}\mathrm{d}x
\int_{-\infty}^\infty
|u'|^2\mathrm{d}x
\ge\frac{(3\alpha+2)^2}{4}
\left(\int_{-\infty}^\infty
|u'|^2|x|^{-\alpha-1}\mathrm{d}r
\right)^2.
\end{align}
Similarly, let $w:=u'$, \eqref{1Dc2} can be rewritten as, for all $w\in C_c^\infty(-\infty,\infty)$,
\begin{align*}
\int_{-\infty}^\infty
|w'|^2|x|^{-2\alpha}\mathrm{d}x
\int_{-\infty}^\infty
|w|^2\mathrm{d}x
\ge\frac{(3\alpha+2)^2}{4}
\left(\int_{-\infty}^\infty
|w|^2|x|^{-\alpha-1}\mathrm{d}x
\right)^2.
\end{align*}
The above inequality and its extremal functions have been proved in \cite[Corollary 1.1-(3),(4)]{Do23}, here we provide a different proof.

Under the assumption $\alpha>-\frac{1}{2}$, let us denote $w(x):=|x|^{2\alpha+1}z(x)$, it follows from H\"{o}lder inequality and integration by parts that
\begin{align*}
\int_{-\infty}^\infty
|w'|^2|x|^{-2\alpha}\mathrm{d}x
\int_{-\infty}^\infty
|w|^2\mathrm{d}x
& =\int_{-\infty}^\infty
|z'|^2|x|^{2\alpha+2}
\mathrm{d}x
\int_{-\infty}^\infty
|z|^2|x|^{4\alpha+2}\mathrm{d}x
\nonumber\\
& \ge \left(\int_{-\infty}^\infty
z'z|x|^{3\alpha+1}x
\mathrm{d}x\right)^2
\nonumber\\
& =\frac{(3\alpha+2)^2}{4}
\left(\int_{-\infty}^\infty
|z|^2|x|^{3\alpha+1}
\mathrm{d}x\right)^2
\nonumber\\
& =\frac{(3\alpha+2)^2}{4}
\left(\int_{-\infty}^\infty |w|^2|x|^{-\alpha-1}
\mathrm{d}x\right)^2,
\end{align*}
and the second equality holds if and only if
\begin{align*}
|z'|^2 |x|^{2\alpha+2}
=C_1|z|^2 |x|^{4\alpha+2}
&\Rightarrow
z(x)=C_2
\exp\left(-c|x|^{\alpha+1}\right)
\\&\Rightarrow
w(x)=C_2|x|^{2\alpha+1}
\exp\left(-c|x|^{\alpha+1}\right)
\ \ \ \ \ \,
(\mbox{by}
\
w(x)=|x|^{2\alpha+1}z(x))
\\&\Rightarrow
u'(x)=C_2|x|^{2\alpha+1}
\exp\left(-c|x|^{\alpha+1}\right)
\qquad\qquad\qquad\ \ \ \,
(\mbox{by}
\
w=u'),
\end{align*}
for some $c,C_1>0$ and $C_2\in\mathbb{R}$. Therefore, the equality in \eqref{1Dc2} holds if and only if $u(x)=a(1+b|x|^{\alpha+1})
\exp\left(-b|x|^{\alpha+1}\right)$ with $a\in \mathbb{R}$ and $b>0$.

Thus, the proof of the case $N=1$ of Theorem \ref{thmwsupsc} is completed.
\end{proof}

Now, inspired by the argument of Cazacu, Flynn and Lam in \cite[Lemma 3.4]{Cazacu22}, we establish the following lemma which will be applied to compute the sharp constant of \eqref{wsUP}.

\begin{lem}\label{lemtrw}
Let $N\geq 2$ and $\alpha>-1$ satisfy $N\geq 5\alpha+5$, then
\begin{equation*}
\inf_{k\in\mathbb{N}\cup\{0\}}
\frac{1}
{\left[1+\frac{4(\alpha+1)k}
{(N+2k-\alpha-3)^2}\right]^2}
\left(\frac{N+2k+3\alpha+1}{2}
\right)^2
=\frac{(N+3\alpha+1)^2}{4}.
\end{equation*}
\end{lem}

\begin{proof}[\rm\textbf{Proof}]
Note that
\begin{align*}
&\frac{1}
{\left[1+\frac{4(\alpha+1)k}
{(N+2k-\alpha-3)^2}\right]^2}
\left(\frac{N+2k+3\alpha+1}{2}
\right)^2
\\&\quad
=\frac{(N+2k-\alpha-3)^4
(N+2k+3\alpha+1)^2}
{4\left[(N+2k-\alpha-3)^2
+4(\alpha+1)k\right]^2}
:=\mathcal{K}(N,\alpha,k).
\end{align*}

$\bullet$ \emph{Step 1: For $N\ge2$ and $\alpha>-1$ satisfying $N\ge 5\alpha+5$, we will verify that}
\[
\inf_{k\in\mathbb{N}}
\mathcal{K}(N,\alpha,k)
=\mathcal{K}(N,\alpha,1).
\]

Under the assumption $N\ge2$ and $\alpha>-1$ satisfying $N\ge5\alpha+5$, we first show that
\begin{align*}
\mathcal{F}(x)
:=\frac{(N+x-\alpha-3)^4
(N+x+3\alpha+1)^2}
{4\left[(N+x-\alpha-3)^2
+2(\alpha+1)x\right]^2}
\end{align*}
is non-decreasing for $x\ge2$. In fact, let $t:=N+x-\alpha-3\ge 4\alpha+4>0$ (due to  $N\ge5\alpha+5$, $x\ge2$ and  $\alpha>-1$), consider the function $\mathcal{G}:[2,\infty)
\times[4\alpha+4,\infty)$:
\begin{align*}
\mathcal{G}(x,t)
=\frac{t^4(t+4\alpha+4)^2}
{4\left[t^2+2(\alpha+1)x\right]^2}.
\end{align*}
The direct computation yields
\begin{align*}
\frac{\partial\mathcal{G}}
{\partial x}(x,t)
&=-\frac{(\alpha+1)t^4
(t+4\alpha+4)^2}
{\left[t^2+2(\alpha+1)x\right]^3},
\\\frac{\partial\mathcal{G}}
{\partial t}(x,t)
&=\frac{(t+4\alpha+4)t^3
\left[t^3+2(\alpha+1)x
(3t+8\alpha+8)\right]}
{2\left[t^2+2(\alpha+1)x\right]^3}.
\end{align*}
Note that $\mathcal{F}(x)=\mathcal{G}(x,t)$ with $t=N+x-\alpha-3$, then we get, for $x\geq 2$ and $t\geq 4\alpha+4>0$,
\begin{align*}
\mathcal{F}'(x)&=
\frac{\partial\mathcal{G}}
{\partial x}(x,t)
+\frac{\partial\mathcal{G}}
{\partial t}(x,t)
\\& =-\frac{(\alpha+1)t^4(t+4\alpha+4)^2}
{\left[t^2+2(\alpha+1)x\right]^3}
+\frac{(t+4\alpha+4)t^3
\left[t^3+2(\alpha+1)x
(3t+8\alpha+8)\right]}
{2\left[t^2
+2(\alpha+1)x\right]^3}
\\&=\frac{(t+4\alpha+4)t^3
\left[t^3+2(\alpha+1)x
(3t+8\alpha+8)\right]
-2(\alpha+1)t^4(t+4\alpha+4)^2}
{2\left[t^2+2(\alpha+1)x\right]^3}
\\&\geq
\frac{(t+4\alpha+4)t^6
-2(\alpha+1)t^4(t+4\alpha+4)^2}
{2\left[t^2+2(\alpha+1)x\right]^3}
\\&=\frac{(t+4\alpha+4)t^4
[t^2-2(\alpha+1)(t+4\alpha+4)]}
{2\left[t^2+2(\alpha+1)x\right]^3}
\\&=\frac{(t+4\alpha+4)t^4
[t+2(\alpha+1)][t-4(\alpha+1)]}
{2\left[t^2+2(\alpha+1)x\right]^3}
\\&\geq 0,
\end{align*}
which means that $\mathcal{F}(x)$ is non-decreasing for $x\ge2$, and then $\mathcal{F}(x)
\ge\mathcal{F}(2)$ for all $x\ge2$. Thus, for all $k\in\mathbb{N}$,
\[
\mathcal{K}(N,\alpha,k)
=\mathcal{F}(2k)
\ge\mathcal{F}(2)
=\mathcal{K}(N,\alpha,1),
\]
that is,
\begin{equation*}
\inf_{k\in\mathbb{N}}
\mathcal{K}(N,\alpha,k)
=\mathcal{K}(N,\alpha,1).
\end{equation*}
In view of this, the proof of \emph{Step 1} is completed.

From \emph{Step 1},
\begin{align}\label{8.11-1}
\inf_{k\in\mathbb{N}\cup\{0\}}
\mathcal{K}(N,\alpha,k)
&=\min\left\{\mathcal{K}(N,\alpha,0),
\mathcal{K}(N,\alpha,1)\right\}.
\end{align}

$\bullet$ \emph{Step 2: For $N\geq 2$ and $\alpha>-1$ satisfying $N\geq 5\alpha+5$, we will show that}
\[
\mathcal{K}(N,\alpha,1)
\ge\mathcal{K}(N,\alpha,0).
\]

The above inequality is equivalent to
\begin{align}\label{8.11-2}
\frac{(N-\alpha-1)^4(N+3\alpha+3)^2}
{4\left[(N-\alpha-1)^2
+4(\alpha+1)\right]^2}
\ge\frac{(N+3\alpha+1)^2}{4}.
\end{align}
Let $y:=N-\alpha-1\geq 4\alpha+4>0$ (due to $N\ge5\alpha+5$ and $\alpha>-1$), then we can rewrite \eqref{8.11-2} as
\begin{equation}\label{8.16-3}
y^4(y+4\alpha+4)^2\geq (y+4\alpha+2)^2(y^2+4\alpha+4)^2.
\end{equation}
Using $N\ge2$, $\alpha>-1$ and $y=N-\alpha-1>0$ again, we see that
\begin{align*}
\begin{cases}
y^2>0,\\
y+4\alpha+4=N+3\alpha+3>0,\\
y+4\alpha+2=N+3\alpha+1>0,\\
y^2+4\alpha+4>0,
\end{cases}
\end{align*}
then \eqref{8.16-3} reduces into
\begin{equation}\label{8.10-1}
y^2\ge2(\alpha+1)y
+4(\alpha+1)(2a+1).
\end{equation}
In order to complete the proof of \emph{Step 2}, it remains to verify that \eqref{8.10-1} holds.

Indeed, by $y\geq 4\alpha+4$ and $\alpha>-1$, then
\begin{align*}
y^2-2(\alpha+1)y
=[y-(\alpha+1)]^2-(\alpha+1)^2
\geq 8(\alpha+1)^2\geq 4(\alpha+1)(2\alpha+1),
\end{align*}
which means that \eqref{8.10-1} holds. This finishes the proof of \emph{Step 2}.

$\bullet$ \emph{Step 3: Conclusion.}

Combining \eqref{8.11-1} and \emph{Step 2},
\begin{equation*}
\inf_{k\in\mathbb{N}\cup\{0\}}
\mathcal{K}(N,\alpha,k)
=\mathcal{K}(N,\alpha,0)
=\frac{(N+3\alpha+1)^2}{4},
\end{equation*}
as our desired estimate. The proof is thereby completed.
\end{proof}

The spherical harmonics decomposition (see \eqref{8.13-1} above) allows us to follow from \cite[(2.9)$-$(2.11)]{Duong21} to obtain the following identities.

\begin{lem}
[{\!\rm{\!\cite[(2.9)--(2.11)]{Duong21}}}]
\label{lemhduukd}
Let $N\geq 2$. For each $u\in C^\infty_c(\mathbb{R}^N):$
\begin{align*}
\int_{\mathbb{R}^N}\frac{|\Delta u|^2}{|x|^{2\alpha}}\mathrm{d}x
& = \sum^\infty_{k=0}
\bigg[\int_0^\infty |v_k''|^2r^{N+2k-2\alpha-1}
\mathrm{d}r
+(2\alpha+1)(N+2k-1)\int_0^\infty |v_k'|^2r^{N+2k-2\alpha-3}
\mathrm{d}r\bigg],
\\ \int_{\mathbb{R}^N}\left|\nabla u\right|^2\mathrm{d}x
& = \sum^\infty_{k=0}\int_0^\infty |v_k'|^2r^{N+2k-1}\mathrm{d}r,
\\ \int_{\mathbb{R}^N}
\frac{\left|\nabla u\right|^2}{|x|^{\alpha+1}}
\mathrm{d}x
&=\sum^\infty_{k=0}
\left[\int_0^\infty |v_k'|^2r^{N+2k-\alpha-2}
\mathrm{d}r
+(\alpha+1)k\int_0^\infty |v_k|^2r^{N+2k-\alpha-4}
\mathrm{d}r
\right].
\end{align*}
\end{lem}

Now, based on the above preparations, we follow the approach of Cazacu, Flynn and Lam \cite[Proof of Theorem 2.3]{Cazacu22} to prove Theorem \ref{thmwsupsc} when $N\ge2$.

\begin{proof}
[\rm\textbf{Proof of Theorem \ref{thmwsupsc}: the case $\boldsymbol{N\ge2}$}]

The proof follows in two steps as follows.

$\bullet$ \emph{Step 1: We will prove that the weighted second order HyUP \eqref{wsUP} holds}.

Let $u\in C^\infty_c(\mathbb{R}^N)$, using  Lemma \ref{lemhduukd}, the  inequality \eqref{wsUP} is equivalent to
\begin{align}\label{8.11-3}
&\sum^\infty_{k=0}
\left[\int_0^\infty |v_k''|^2r^{N+2k-2\alpha-1}
\mathrm{d}r
+(2\alpha+1)(N+2k-1)
\int_0^\infty |v_k'|^2r^{N+2k-2\alpha-3}
\mathrm{d}r
\right]
\nonumber\\
&\qquad \times\sum^\infty_{k=0}
\int_0^\infty |v_k'|^2r^{N+2k-1}\mathrm{d}r
\nonumber\\
&\quad\ge C(N,\alpha)
\left\{\sum^\infty_{k=0}
\left[\int_0^\infty |v_k'|^2r^{N+2k-\alpha-2}
\mathrm{d}r
+(\alpha+1)k\int_0^\infty |v_k|^2r^{N+2k-\alpha-4}
\mathrm{d}r
\right]\right\}^2.
\end{align}
By Cauchy-Schwarz inequality, in order to verify \eqref{8.11-3}, it suffices to show that, for all $k\in\mathbb{N}\cup\{0\}$,
\begin{align}\label{8.11-4}
&\left[\int_0^\infty |v_k''|^2r^{N+2k-2\alpha-1}
\mathrm{d}r
+(2\alpha+1)(N+2k-1)\int_0^\infty |v_k'|^2r^{N+2k-2\alpha-3}
\mathrm{d}r
\right]\int_0^\infty |v_k'|^2r^{N+2k-1}\mathrm{d}r
\nonumber\\
&\quad \ge C(N,\alpha)
\left[\int_0^\infty |v_k'|^2r^{N+2k-\alpha-2}
\mathrm{d}r
+(\alpha+1)k\int_0^\infty |v_k|^2r^{N+2k-\alpha-4}
\mathrm{d}r
\right]^2.
\end{align}
For $k\in\mathbb{N}$, applying integration by parts and H\"{o}lder inequality, we arrive at
\begin{align*}
\int_0^\infty |v_k|^2r^{N+2k-\alpha-4}
\mathrm{d}r
&=\frac{1}{N+2k-\alpha-3}
\int_0^\infty |v_k|^2
\left(r^{N+2k-\alpha-3}\right)'
\mathrm{d}r
\\&=\frac{-2}{N+2k-\alpha-3}
\int_0^\infty v_kv_k'
r^{N+2k-\alpha-3}
\mathrm{d}r
\\&\le\frac{2}{N+2k-\alpha-3}
\left(\int_0^\infty \!|v_k'|^2r^{N+2k-\alpha-2}
\mathrm{d}r\!\right)^{\frac{1}{2}}
\left(\int_0^\infty \!|v_k|^2r^{N+2k-\alpha-4}
\mathrm{d}r\!\right)^{\frac{1}{2}},
\end{align*}
which is equivalent to
\[
\int_0^\infty |v_k|^2r^{N+2k-\alpha-4}
\mathrm{d}r
\le\frac{4}{(N+2k-\alpha-3)^2}
\int_0^\infty |v_k'|^2r^{N+2k-\alpha-2}
\mathrm{d}r,
\]
this gives
\begin{align*}
&\int_0^\infty |v_k'|^2r^{N+2k-\alpha-2}
\mathrm{d}r
+(\alpha+1)k\int_0^\infty |v_k|^2r^{N+2k-\alpha-4}
\mathrm{d}r
\\&\quad\le
\left[1+\frac{4(\alpha+1)k}
{(N+2k-\alpha-3)^2}\right]
\int_0^\infty |v_k'|^2r^{N+2k-\alpha-2}
\mathrm{d}r.
\end{align*}
Thus, to verify \eqref{8.11-4},  it is enough to prove that, for any $k\in\mathbb{N}\cup\{0\}$,
\begin{align*}
&\left[\int_0^\infty |v_k''|^2r^{N+2k-2\alpha-1}
\mathrm{d}r
+(2\alpha+1)(N+2k-1)\int_0^\infty |v_k'|^2r^{N+2k-2\alpha-3}
\mathrm{d}r
\right]\int_0^\infty |v_k'|^2r^{N+2k-1}\mathrm{d}r
\nonumber\\
&\quad \ge C(N,\alpha)
\left[1+\frac{4(\alpha+1)k}
{(N+2k-\alpha-3)^2}\right]^2
\left(\int_0^\infty |v_k'|^2r^{N+2k-\alpha-2}
\mathrm{d}r\right)^2.
\end{align*}
In view of this, we try to get the sharp constant $D(N,\alpha,k)$ in the following inequality:
\begin{align}\label{1dni}
&
\left[\int_0^\infty |v_k''|^2r^{N+2k-2\alpha-1}
\mathrm{d}r
+(2\alpha+1)(N+2k-1)\int_0^\infty |v_k'|^2r^{N+2k-2\alpha-3}
\mathrm{d}r
\right]\int_0^\infty |v_k'|^2r^{N+2k-1}\mathrm{d}r
\nonumber\\
&\quad \ge D(N,\alpha,k)\left(\int_0^\infty |v_k'|^2r^{N+2k-\alpha-2}
\mathrm{d}r\right)^2,
\end{align}
for any $k\in\mathbb{N}\cup\{0\}$. Then
\begin{equation*}
C(N,\alpha)
\geq\inf_{k\in\mathbb{N}\cup\{0\}}
\frac{D(N,\alpha,k)}
{\left[1+\frac{4(\alpha+1)k}
{(N+2k-\alpha-3)^2}\right]^2}.
\end{equation*}
Let $w:=v_k'$, then \eqref{1dni} becomes, for any $k\in\mathbb{N}\cup\{0\}$,
\begin{align*}
&\left[\int_0^\infty |w'|^2r^{N+2k-2\alpha-1}\mathrm{d}r
+(2\alpha+1)(N+2k-1)\int_0^\infty |w|^2r^{N+2k-2\alpha-3}\mathrm{d}r
\right]\int_0^\infty |w|^2r^{N+2k-1}\mathrm{d}r
\nonumber\\
&\quad \ge D(N,\alpha,k)\left(\int_0^\infty |w|^2r^{N+2k-\alpha-2}
\mathrm{d}r\right)^2.
\end{align*}
Fix $0<\varepsilon\ll 1$. Let $w(r):=r^{2\alpha+1}z(r)$ for  $r\in[\varepsilon,\infty)$. Then applying integration by parts, we see that
\begin{align*}
&\int_\varepsilon^\infty |w'|^2r^{N+2k-2\alpha-1}
\mathrm{d}r
+(2\alpha+1)(N+2k-1)
\int_\varepsilon^\infty |w|^2r^{N+2k-2\alpha-3}
\mathrm{d}r
\\&\quad=
\int_\varepsilon^\infty \left|(2\alpha+1)r^{2\alpha}z
+r^{2\alpha+1}z'\right|^2
r^{N+2k-2\alpha-1}
\mathrm{d}r
\nonumber\\&\qquad
+(2\alpha+1)(N+2k-1)
\int_\varepsilon^\infty |z|^2r^{N+2k+2\alpha-1}
\mathrm{d}r
\\&\quad=
(2\alpha+1)^2
\int_\varepsilon^\infty \left|z\right|^2
r^{N+2k+2\alpha-1}
\mathrm{d}r
+\int_\varepsilon^\infty
\left|z'\right|^2
r^{N+2k+2\alpha+1}
\mathrm{d}r
\nonumber\\&\qquad
+2(2\alpha+1)
\int_\varepsilon^\infty
zz'r^{N+2k+2\alpha}
\mathrm{d}r
+(2\alpha+1)(N+2k-1)
\int_\varepsilon^\infty |z|^2r^{N+2k+2\alpha-1}
\mathrm{d}r
\\&\quad=
(2\alpha+1)^2
\int_\varepsilon^\infty \left|z\right|^2
r^{N+2k+2\alpha-1}
\mathrm{d}r
+\int_\varepsilon^\infty
\left|z'\right|^2
r^{N+2k+2\alpha+1}
\mathrm{d}r
\nonumber\\&\qquad
-(2\alpha+1)(N+2k+2\alpha)
\int_\varepsilon^\infty
|z|^2
r^{N+2k+2\alpha-1}
\mathrm{d}r+o_\varepsilon(1)
\nonumber\\&\qquad
+(2\alpha+1)(N+2k-1)
\int_\varepsilon^\infty |z|^2r^{N+2k+2\alpha-1}
\mathrm{d}r
\\&\quad=
\int_\varepsilon^\infty |z'|^2r^{N+2k+2\alpha+1}
\mathrm{d}r
+o_\varepsilon(1).
\end{align*}
Hence, using integrating by parts and H\"{o}lder inequality, we get
\begin{align}\label{1dnibk}
&
\left[\int_\varepsilon^\infty |w'|^2r^{N+2k-2\alpha-1}
\mathrm{d}r
+(2\alpha+1)(N+2k-1)
\int_\varepsilon^\infty |w|^2r^{N+2k-2\alpha-3}
\mathrm{d}r
\right]\int_\varepsilon^\infty |w|^2r^{N+2k-1}\mathrm{d}r
\nonumber\\
&\qquad =\left[\int_\varepsilon^\infty |z'|^2r^{N+2k+2\alpha+1}
\mathrm{d}r
+o_\varepsilon(1)\right]
\int_\varepsilon^\infty |z|^2r^{N+2k+4\alpha+1}
\mathrm{d}r
\nonumber\\
&\qquad =\int_\varepsilon^\infty |z'|^2r^{N+2k+2\alpha+1}
\mathrm{d}r
\int_\varepsilon^\infty |z|^2r^{N+2k+4\alpha+1}
\mathrm{d}r
+o_\varepsilon(1)
\nonumber\\
&\qquad \ge \left(\int_\varepsilon^\infty zz'r^{N+2k+3\alpha+1}
\mathrm{d}r\right)^2
+o_\varepsilon(1)
\nonumber\\
&\qquad =\frac{1}{4}
\left[\int_\varepsilon^\infty
\left(|z(r)|^2\right)'
r^{N+2k+3\alpha+1}
\mathrm{d}r\right]^2
+o_\varepsilon(1)
\nonumber\\
&\qquad =\frac{(N+2k+3\alpha+1)^2}{4}
\left(\int_\varepsilon^\infty
|z(r)|^2 r^{N+2k+3\alpha}
\mathrm{d}r\right)^2
+o_\varepsilon(1)
\nonumber\\
&\qquad =\frac{(N+2k+3\alpha+1)^2}{4}
\left(\int_\varepsilon^\infty |w|^2r^{N+2k-\alpha-2}
\mathrm{d}r\right)^2
+o_\varepsilon(1).
\end{align}
Let $\varepsilon\to 0$, we obtain, for any $k\in\mathbb{N}\cup\{0\}$,
\begin{align*}
&\left[\int_0^\infty |w'|^2r^{N+2k-2\alpha-1}
\mathrm{d}r
+(2\alpha+1)(N+2k-1)\int_0^\infty |w|^2r^{N+2k-2\alpha-3}
\mathrm{d}r
\right]\int_0^\infty |w|^2r^{N+2k-1}\mathrm{d}r
\nonumber\\
&\quad \ge \frac{(N+2k+3\alpha+1)^2}{4}
\left(\int_0^\infty |w|^2r^{N+2k-\alpha-2}
\mathrm{d}r\right)^2.
\end{align*}
Thus, under the assumptions  $N\geq 2$ and $\alpha>-1$ satisfying $N\geq 5\alpha+5$, we deduce from Lemma \ref{lemtrw} that
\begin{align*}
C(N,\alpha)
\geq\inf_{k\in\mathbb{N}\cup\{0\}}
\frac{1}
{\left[1+\frac{4(\alpha+1)k}
{(N+2k-\alpha-3)^2}\right]^2}
\frac{(N+2k+3\alpha+1)^2}{4}
=\frac{(N+3\alpha+1)^2}{4}.
\end{align*}
Therefore, by standard density argument, for all $u\in H^2_{\alpha,0}(\mathbb{R}^N)$, the desired result \eqref{wsUP} follows.

$\bullet$ \emph{Step 2: We will verify the attainability of the sharp constant $\frac{(N+3\alpha+1)^2}
{4}$} of \eqref{wsUP}.

From Lemma \ref{lemtrw}, we see that
\begin{align*}
&\inf_{k\in\mathbb{N}\cup\{0\}}
\frac{1}
{\left[1+\frac{4(\alpha+1)k}
{(N+2k-\alpha-3)^2}\right]^2}
\left(\frac{N+2k+3\alpha+1}{2}
\right)^2
\\&\quad
=\inf_{k\in\mathbb{N}\cup\{0\}}
\frac{1}
{\left[1+\frac{4(\alpha+1)k}
{(N+2k-\alpha-3)^2}\right]^2}
\left(\frac{N+2k+3\alpha+1}{2}
\right)^2\bigg|_{k=0}
=\frac{(N+3\alpha+1)^2}{4},
\end{align*}
then if the extremal function of \eqref{wsUP} exists, it must be radial (according to $k=0$). Moreover, if $u(|x|)$ is the extremal function of \eqref{wsUP}, then $w(r):=u'(r)$ also must be the extremal function of
\begin{align}\label{wuppks}
&\left[
\int^\infty_0\left|w'\right|^2
r^{N-2\alpha-1}\mathrm{d}r
+(2\alpha+1)(N-1)
\int^\infty_0\left|w\right|^2
r^{N-2\alpha-3}\mathrm{d}r\right]
\int^\infty_0\left|w\right|^2
r^{N-1}\mathrm{d}r
\nonumber
\\
&\quad \geq \frac{(N+3\alpha+1)^2}{4}
\left(\int^\infty_0\left|w\right|^2
r^{N-\alpha-2}\mathrm{d}r\right)^2.
\end{align}
By \eqref{1dnibk} with $k=0$, we know that the equality in \eqref{wuppks} holds when applying H\"{o}lder inequality to \eqref{1dnibk} if and only if
\begin{align*}
|z'|^2 r^{N+2\alpha+1}
=C_1|z|^2 r^{N+4\alpha+1}
&\Rightarrow
z(r)=C_2
\exp\left(-cr^{\alpha+1}\right)
\\&\Rightarrow
w(r)=C_2r^{2\alpha+1}
\exp\left(-cr^{\alpha+1}\right)
\ \ \ \ \ \ \ \ \ \,
(\mbox{by}
\
w(r)=r^{2\alpha+1}z(r))
\\&\Rightarrow
u'(r)=C_2r^{2\alpha+1}
\exp\left(-cr^{\alpha+1}\right)
\qquad\qquad\ \ \ \,
(\mbox{by}
\
w(r)=u'(r)),
\end{align*}
for some $c,C_1>0$ and $C_2\in\mathbb{R}$.
Therefore, the extremal function of \eqref{wsUP} must be of the form $\bar{u}(x)=a(1+b|x|^{\alpha+1})
\exp\left(-b |x|^{\alpha+1}\right)$ with $a\in \mathbb{R}$ and $b>0$.

Indeed, direct computations yield
\begin{align*}
\int_{\mathbb{R}^N}\frac{|\Delta \bar{u}|^2}{|x|^{2\alpha}}
\mathrm{d}x
&=\left|\mathbb{S}^{N-1}\right|
\int^\infty_0
\left|\bar{u}''+\frac{N-1}{r}
\bar{u}'\right|^2
r^{N-2\alpha-1}\mathrm{d}r
\\&=\left|\mathbb{S}^{N-1}\right|
a^2b^4(\alpha+1)^2
\int^\infty_0
\left|(N+2\alpha)-b(\alpha+1)
r^{\alpha+1}\right|^2
\exp\left(-2br^{\alpha+1}\right)
r^{N+2\alpha-1}\mathrm{d}r
\\&=\left|\mathbb{S}^{N-1}\right|
a^2b^4(\alpha+1)^2
\bigg[(N+2\alpha)^2
\int^\infty_0
\exp\left(-2br^{\alpha+1}\right)
r^{N+2\alpha-1}\mathrm{d}r
\\&\quad
-2(N+2\alpha)b(\alpha+1)
\int^\infty_0
\exp\left(-2br^{\alpha+1}\right)
r^{N+3\alpha}
\mathrm{d}r
\nonumber\\&\quad
+b^2(\alpha+1)^2
\int^\infty_0
\exp\left(-2br^{\alpha+1}\right)
r^{N+4\alpha+1}\mathrm{d}r\bigg].
\end{align*}
Next, we change the variable: $s=2br^{\alpha+1}$, then
\begin{align*}
\int^\infty_0
\exp\left(-2br^{\alpha+1}\right)
r^{N+2\alpha-1}
\mathrm{d}r
&=\frac{1}{(\alpha+1)
(2b)^{\frac{N+2\alpha}{\alpha+1}}}
\int^\infty_0
e^{-s}
s^{\frac{N+2\alpha}{\alpha+1}-1}
\mathrm{d}s
\\&=\frac{1}{(\alpha+1)
(2b)^{\frac{N+2\alpha}{\alpha+1}}}
\Gamma\left(\frac{N+2\alpha}
{\alpha+1}\right),
\\
\int^\infty_0
\exp\left(-2br^{\alpha+1}\right)
r^{N+3\alpha}\mathrm{d}r
&=\frac{1}{(\alpha+1)
(2b)^{\frac{N+2\alpha}{\alpha+1}+1}}
\int^\infty_0
e^{-s}s^{\frac{N+2\alpha}{\alpha+1}}
\mathrm{d}s
\\&=\frac{1}{(\alpha+1)
(2b)^{\frac{N+2\alpha}{\alpha+1}+1}}
\Gamma\left(\frac{N+2\alpha}
{\alpha+1}+1\right),
\\
\int^\infty_0
\exp\left(-2br^{\alpha+1}\right)
r^{N+4\alpha+1}\mathrm{d}r
&=\frac{1}{(\alpha+1)
(2b)^{\frac{N+2\alpha}{\alpha+1}+2}}
\int^\infty_0
e^{-s}s^{\frac{N+2\alpha}{\alpha+1}+1}
\mathrm{d}s
\\&=\frac{1}{(\alpha+1)
(2b)^{\frac{N+2\alpha}{\alpha+1}+2}}
\Gamma\left(\frac{N+2\alpha}
{\alpha+1}+2\right).
\end{align*}
Hence, using the property of the Gamma function: $\Gamma(t+1)=t\Gamma(t)$ for $t>0$, there holds
\begin{align*}
\int_{\mathbb{R}^N}\frac{|\Delta \bar{u}|^2}{|x|^{2\alpha}}
\mathrm{d}x
=\frac{\left|\mathbb{S}^{N-1}\right|
a^2b^4(\alpha+1)}
{4(2b)^{\frac{N+2\alpha}{\alpha+1}}}
(N+3\alpha+1)(N+2\alpha)
\Gamma\left(\frac{N+2\alpha}
{\alpha+1}\right).
\end{align*}
Also,
\begin{align*}
\int_{\mathbb{R}^N}
|\nabla \bar{u}|^2\mathrm{d}x
&=\left|\mathbb{S}^{N-1}\right|
\int^\infty_0
\left|\bar{u}'\right|^2
r^{N-1}\mathrm{d}r
\\&=\left|\mathbb{S}^{N-1}\right|
a^2b^4(\alpha+1)^2
\int^\infty_0
\exp\left(-2br^{\alpha+1}\right)
r^{N+4\alpha+1}\mathrm{d}r
\\&=
\frac{\left|\mathbb{S}^{N-1}\right|
a^2b^4(\alpha+1)}
{(2b)^{\frac{N+2\alpha}{\alpha+1}+2}}
\Gamma\left(\frac{N+2\alpha}
{\alpha+1}+2\right),
\end{align*}
and
\begin{align*}
\int_{\mathbb{R}^N}
\frac{|\nabla \bar{u}|^2}{|x|^{\alpha+1}}
\mathrm{d}x
&=\left|\mathbb{S}^{N-1}\right|
\int^\infty_0
\left|\bar{u}'\right|^2
r^{N-\alpha-2}\mathrm{d}r
\\&=\left|\mathbb{S}^{N-1}\right|
a^2b^4(\alpha+1)^2
\int^\infty_0
\exp\left(-2br^{\alpha+1}\right)
r^{N+3\alpha}\mathrm{d}r
\\&=
\frac{\left|\mathbb{S}^{N-1}\right|
a^2b^4(\alpha+1)}
{(2b)^{\frac{N+2\alpha}{\alpha+1}+1}}
\Gamma\left(\frac{N+2\alpha}
{\alpha+1}+1\right).
\end{align*}
Therefore,
\begin{align*}
\dfrac{\int_{\mathbb{R}^N}
\frac{|\Delta \bar{u}|^2}{|x|^{2\alpha}}
\mathrm{d}x
\int_{\mathbb{R}^N}
\left|\nabla \bar{u}\right|^2 \mathrm{d}x}
{\left(\int_{\mathbb{R}^N}
\frac{\left|\nabla \bar{u}\right|^2}{|x|^{\alpha+1}}
\mathrm{d}x\right)^2}
& =\frac{\frac{(N+3\alpha+1)
(N+2\alpha)
\Gamma\left(\frac{N+2\alpha}
{\alpha+1}\right)}
{4(2b)^{\frac{N+2\alpha}{\alpha+1}}}
\frac{\Gamma
\left(\frac{N+2\alpha}
{\alpha+1}+2\right)}
{(2b)^{\frac{N+2\alpha}{\alpha+1}+2}}}
{\left[\frac{\Gamma
\left(\frac{N+2\alpha}
{\alpha+1}+1\right)}
{(2b)^{\frac{N+2\alpha}{\alpha+1}+1}}
\right]^2}
\\& =\frac{(N+3\alpha+1)
(N+2\alpha)
\left(\frac{N+2\alpha}
{\alpha+1}+1\right)
\frac{N+2\alpha}{\alpha+1}}
{4\left(\frac{N+2\alpha}
{\alpha+1}\right)^2}
\\&=\frac{(N+3\alpha+1)^2}{4}.
\end{align*}
Now, the proof of Theorem \ref{thmwsupsc} is thereby completed.
\end{proof}

\section*{Declarations}

\subsection*{Funding}
\noindent This paper was supported by the National Natural Science Foundation of China (No. 12371120).

\subsection*{Data availability statement}
\noindent No data was used for the research described in the article.

\subsection*{Conflict of interest}
\noindent The authors declare no conflict of interest.

\end{document}